OpenStudy (anonymous):

Is this equal? (X - 1)(x^4 - 5x^2 + 4)= (x^5 - 5x^3 + 4x) or is it: (X - 1)(-x^4 + 5x^2 - 4)= (x^5 - 5x^3 + 4x)

OpenStudy (anonymous):

The second polynomial is the original . I need to know if I factored out x-1 correctly.

OpenStudy (campbell_st):

split the binomial factor \[(x -1)(x^4 - 5x^2 + 4) = x \times(x^4 - 5x^2 + 4) - 1 \times (x^4 - 5x^2 + 4)\] distribute and collect like terms

OpenStudy (campbell_st):

well it hasn't been factored correctly, as there is only 1 common factor